OpenStudy (anonymous):

using radical notation :

OpenStudy (anonymous):

\[\sqrt[3]{54m ^{9}n ^{5}}\] simplify

OpenStudy (anonymous):

alright let's take it a step at a time :) 54 = 9 x 6 = 3x3x3x2, so : \[54 =( 3^3 )(2)\] \[m^9 = (m^3)(m^3)(m^3)\] so you can write it down as : \[m^9 = (m^3)^3\] and \[n^5 = (n^3)(n^2)\] now let's write these down under the cubic root ^_^ : \[\sqrt[3]{(3)^3(2)(m^3)^3(n)^3(n)^2}\] now we can take the ones that have a cubic power out , so the answer is : \[= 3m^3n^3\sqrt[3]{2n^2}\] understand it ? ^_^

OpenStudy (anonymous):

sorry it should be :\[3m^3n \sqrt[3]{2n^2}\] :)

OpenStudy (anonymous):

what happend to the extra 3 that was after (m^3) ? cancelled out?

OpenStudy (anonymous):

you have a cubic root right? so m^3's power which was also another(3) was cancelled ^_^

OpenStudy (anonymous):

lol, yeah :)